OpenStudy (anonymous):

Find all solutions if 0° ≤ θ < 360°. Verify your answer graphically. (Enter your answers from smallest to largest.) cos(3theta)=1

OpenStudy (aum):

\[ \cos(3\theta)=1 \\ \text{From unit circle, cosine is 1 when theta = 0} \\ \text{Also cosine has a period of 2pi. Therefore, 0, 0+2pi, 0+4pi, ... are solutions.} \\ 3\theta = 2n\pi ~~ \text{where }n = 0, 1, 2, 3, .... \]

OpenStudy (aum):

In degrees, 0, 0 +360, 0 + 720, .... are all solutions. \[ 3\theta = 0, ~~360, ~~720, ~~1080,~~ 1440, ....\\ \theta =~~0, ~~120, ~~240, ~~~~360 \]

OpenStudy (anonymous):

wait zero is an option?

OpenStudy (anonymous):

are u in the red cross? or are u swiss.. ur image is distracting

OpenStudy (aum):

Neither. Just an image off the net.

OpenStudy (anonymous):

hahah ok...

OpenStudy (aum):

So 0, 120, 240 and 360 are the solutions.

OpenStudy (anonymous):

wouldnt 240 be an option?

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

yay thank you! got it right... i had the right answers but i forgot zero

OpenStudy (anonymous):

can you help me on another one?

OpenStudy (aum):

yw.

OpenStudy (aum):

go ahead.

OpenStudy (anonymous):

cos(2theta)= - 1/2 so i got... -60... so 300 and 60 because its positive... right? so i thought it either has to be in q1 or q4 ... so i got 150, 30 etc... well i got it wrong...

OpenStudy (aum):

Do they specify the domain of theta like they did in the previous problem?

OpenStudy (anonymous):

360

OpenStudy (aum):

The cosine function is negative in the second and third quadrant. From unit circle, cosine is -1/2 when angle = 120 and 240. cosine had a periodicity of 360 degrees. So 120, 240, 120+360=480, 240+360=600 all give -1/2 for cosine. 2(theta) = 120, 240, 480, 600, ... theta = 60, 120, 240, 300.

OpenStudy (anonymous):

so subtract 60 from 180 and add to 180? ?

OpenStudy (aum):

An acronym to remember is ASTC or All Students Take Calculus. A: All trig functions are positive in the first quadrant. S: Sine & Csc are positive in the second quadrant. T: Tan & Cot are positive in the third quadrant. C: Cosine and Sec are positive in the fourth quadrant.

OpenStudy (anonymous):

ohh thanks thats pretty cool

OpenStudy (aum):

Here cosine is negative half. So it must be in the second and third quadrant. Cosine is positive half when angle = 60 degrees. So the reference angle is 60 degrees to get half. To get negative half just use the reference angle of 60 degrees to arrive at angles in q@ and Q3. (Remember reference angle is the angle with the x-axis). So 180 - 60 = 120 180 + 60 = 240 So cos(120) = cos(240) = -1/2. After that make use of the periodicity of cosine which is 360 degrees. Keep adding as many multiples of 360 degrees to 120 and 240 to arrive at other solutions. So cos(2*theta) = -1/2 2*theta = 120, 240, 120+360=480, 240+360=600, .... theta = 60, 120, 240, 300.

OpenStudy (aum):

To get negative half just use the reference angle of 60 degrees to arrive at angles in Q2 and Q3.

OpenStudy (aum):

All trig functions have this property: Measure the angle made with the x-axis. It can be with the positive x-axis if that gives the acute angle. Or it can be with the negative x-axis if that gives the acute angle. |sin(x)| = |sin(180-x)| = |sin(180+x)| = |sin(360-x)|. This is true for all trig functions. The magnitude is the same. They differ only in sign. By remembering ASTC we can figure out the sign.

OpenStudy (aum):



OpenStudy (aum):

|tan(x)| = |tan(180-x)| = |tan(180+x)| = |tan(360-x)| For example, in the first quadrant if you just remembered that tan(45) = 1 and that tan is positive in the first and third quadrant, you can say tan(180+45) = tan(45) = 1 and tan(180-45) = tan(360-45) = -1

OpenStudy (aum):

So knowing the value of the trig function for angle x in the first quadrant, you will know the value of the trig functions for the angles: x (Q1), 180-x (Q2), 180+x (Q3), and 360-x (Q4) right away.

OpenStudy (aum):

Example sin(30) = 1/2 Sine is positive in Q1 and Q2 (from ASTC). 30 is Q1, 180-30=150 is Q2, 180+30=210 is Q3, 360-30 = 330 is Q4. sin(30) = 1/2. Therefore, sin(150) = 1/2 sin(210) = -1/2 sin(330) = -1/2

OpenStudy (aum):

Another example, cos(60) = 1/2. Reference angle is 60 degrees. The other angles in Q2, Q3, Q4 from the reference angle of 60 degrees are: Q1: 60 Q2: 180 - 60 =120 Q3: 180 + 60 = 240 Q4: 360 - 60 = 300 Cosine is positive in Q1 and Q4. Therefore, from cos(60) = 1/2 we can conclude: cos(120) = -1/2 cos(240) = -1/2 cos(300) = 1/2

OpenStudy (aum):

I think I may have confused you enough for the day. :)

OpenStudy (anonymous):

i totally get it.. i just got lost with other problems in my hw